Below are given the prices and maturities of zero-coupon bonds of the country X. Each bond has a face value of 1000.

Maturity (years)

Current price

1

952.38

2

942.60

3

928.60

What is the yield on the bond that matures in three years?

Please write an answer in decimals. For example, 12.3% would be 0.123.
Also, round your answer to the third decimal.
Answer #1

yield on bond that matures in 3 years=(face value/price of 3 year bond)^(1/3)-1=(1000/928.60)^(1/3)-1=2.50%